# VALUE: Value-Aware Large Language Model for Query Rewriting via Weighted Trie in Sponsored Search

### v1

@@ Abstract.
In the realm of sponsored search advertising, matching advertisements with the search intent of a user’s query is crucial. Query-to-bidwords(i.e. bidding keywords) rewriting, which involves transforming user queries into keywords for bidding, is a vital technique that has garnered significant attention from both industry and academia. Recently, with the prevalence of large language models (LLMs), generative retrieval methods have proven effective in producing high-relevance rewrites. However, we have identified a significant limitation in existing approaches: While fine-tuning LLMs for specific domains enhances semantic relevance, these models have no perception of the intrinsic value of their generated outputs, such as commercial value. Therefore, after supervised fine-tuning (SFT), a reinforcement learning from human feedback (RLHF) phase is often employed to address this issue. Nevertheless, traditional preference alignment methods often face challenges in aligning fine-grained values and are susceptible to overfitting, which diminishes the effectiveness and quality of the generated results. To address these challenges, we propose VALUE (Value-Aware Large language model for qUery rewriting via wEighted trie), the first framework that ensures the generation of high-value and highly relevant bidwords. Our approach utilizes weighted trie, an innovative modification of the traditional trie data structure. By modulating the LLM’s output probability distribution with value information from the trie during decoding process, we constrain the generation space and guide the trajectory of text production. Our method not only addresses fine-grained value alignment but also effectively reduces the hallucination issues often encountered with LLMs. Offline experiments demonstrate the effectiveness of our method in semantic matching and preference alignment, showing a remarkable improvement in the value attribute by more than fivefold. Online A/B tests further revealed that our Revenue Per Mille (RPM) metric increased by 1.64%. VALUE has been deployed on our advertising system since October 2024 and served the Double Eleven promotions, the biggest shopping carnival in China.
@@
These authors contributed equally to this work and should be considered co-first authors.
This work was done when the first author was an intern at Alimama.
@@ 1. Introduction
Figure 1. System Overview of our proposed VALUE Framework includes three parts: ”SFT and Alignment”, ”Weighted Trie Construction and Weight Momentum Update”, and ”Collaborative Inference”. During the LLM decoding process, we derive a value distribution from a weighted trie to constrain the generation space, and subsequently merge this with another distribution to produce the final output distribution.
[…]
Sponsored search is a crucial element of modern search engines, where advertisers bid on relevant bidwords to display their ads alongside original search results. Upon receiving user’s query, it is essential to match the query with relevant bidwords (Fain and Pedersen, 2006), and subsequently identify suitable ads associated with those bidwords. This process is a critical component in the ad retrieval pipeline, driving the major revenue of search engines (Jansen and Mullen, 2008). The success of sponsored search depends on the search engine’s ability to accurately identify bidwords that closely align with user intent. However, users often face challenges when formulating queries, leading to suboptimal search experiences. These challenges are intensified when users do not use precise or correct terminology, employ synonyms, or mix languages in their search phrases due to varying levels of language proficiency. Additionally, search terms might be misspelled or overly general, complicating the retrieval of relevant ads. For example, a user might search for ”spring fashion,” which is broad and could correspond to bidwords like ”2024 spring collection” or ”women’s spring trends”. Each query reflects different intents but lacks clarity without additional context. As e-commerce platforms continue to grow in both scale and diversity, ensuring accurate and relevant ad retrieval becomes increasingly challenging, necessitating the need for advanced query rewriting and bidword matching techniques to better align user queries with advertiser bidwords.
In query rewriting, two primary paradigms exist: discriminative methods, which reformulate queries by the retrieval of similar terms and generative methods, which utilize language models for direct query-to-bidword transformation. Discriminative methods leverage sparse and dense retrieval techniques to find relevant bidwords. Sparse retrieval techniques, such as BM25 (Robertson et al., 2009), employ exact lexical matching through high-dimensional term-frequency vectors. Their reliance on exact lexical matching leads to semantic brittleness. This rigidity prevents them from recognizing semantic equivalences between different expressions, resulting in significant retrieval failures when faced with semantically similar but lexically divergent queries. Dense retrieval approaches overcome the limitations of sparse methods by utilizing low-dimensional semantic embeddings generated by neural encoders like sBERT (Reimers, 2019). Dense retrieval, while effective in capturing semantic relationships, suffers from significant limitations, particularly the Matthew Effect and long-tail challenges. High-frequency queries dominate model training due to abundant click logs, leading to over-optimized embeddings for head queries while neglecting tail queries with sparse data. Furthermore, dense retrieval models struggle with distribution shifts and require costly strategies like meta-learning or data augmentation to address long-tail issues (Peng et al., 2023).
Recently, the rise of generative retrieval has garnered widespread attention. Wang (Wang et al., 2024) utilizes LLMs for one-step query-to-keyword rewriting, using a trie during the decoding process. Despite improvements over traditional information retrieval (IR) baselines, existing generative approaches using LLMs for query-to-bidword rewriting tasks have notable limitations. A key limitation is their inability to recognize the value of the generated results, which, in our scenario, pertains to the outcomes they can deliver. Models trained solely through supervised fine-tuning on online logs lack awareness of the value of the results. Due to the training methodology and the SFT loss function design, the model treats each query-bidword pair in the training samples equally, failing to learn any preference or value information. Although there is a great deal of work on Reinforcement Learning from Human Feedback (RLHF), such as PPO, DPO, KTO, and ORPO(Ouyang et al., 2022; Rafailov et al., 2024; Ethayarajh et al., 2024; Hong et al., 2024), which can partially mitigate the value perception issues of SFT models, these methods also have significant drawbacks. Firstly, they require substantial computational resources for training. However, the effective Cost Per Mille (eCPM) value of bidwords is a statistical measure that frequently fluctuates. When these fluctuations occur, the previous alignment efforts become obsolete, and the model needs to be retrained. Secondly, these training methods are prone to overfitting (Meng et al., 2024), especially in our scenario. Unlike traditional alignment objectives for LLMs, such as harmlessness and usefulness, which can be achieved using a small dataset, the value of a bidword is completely independent of its textual content. Therefore, the model must be trained individually for each bidword. Our experiment on DPO show that models trained in this manner tend to overfitting. There is an urgent need for a new paradigm to address the challenges above.
Through our observations, we have identified an opportunity in the sponsored search advertising scenario to enhance the value of bidwords without compromising their relevance. The value of a bidword does not necessarily correlate with its literal meaning. Two bidwords might differ by only a symbol or character, yet exhibit substantial differences in value. Therefore, there is an opportunity to increase the value of the top $K$ bidwords while minimally affecting relevance by prioritizing those with the same level of relevance but higher value over others. Thereby, we present VALUE, an innovative reward-guided inference framework that integrates value-awareness into the decoding process via a novel weighted trie structure. We design algorithms for the construction and update of the value of trie node. During decoding process, VALUE obtain reward information from for the next tokens the weighted trie. This information is used to adjust the output distribution, thereby sampling the subsequent token. By this reward-guided search, the framework aims to maximize the reward of the outputs. We used a multi-task fine-tuning strategy to equip the model with essential query rewriting skills. To achieve better value alignment, we designed an optional post-training alignment phase. To evaluate the effectiveness of our proposed method, we conducted offline experiments using historical online logs. Compared to the current online service model, our method demonstrated a 9.3% improvement in relevance. Additionally, online A/B tests showed increases in key profit metrics, with Cost rising by 1.53% and Revenue Per Mille (RPM) increasing by 1.64%. Given these significant improvements, our method has been fully deployed in the online service. Our contributions can be summarized as follows:
@@
    Pioneering Query Rewriting Framework via Collaborative Inference: We present a groundbreaking query rewriting framework that integrates LLMs with weighted trie, uniquely designed to simultaneously optimize for both high value and semantic relevance. This framework represents the first of its kind to effectively balance these two critical aspects.
@@
    Renovation of Trie: We have reinvented the trie data structure, by seamlessly integrating value information directly into its nodes. This advancement is further supported by our proposed algorithms for trie construction and momentum update of node value.
@@
    Post-Training Alignment for Fine-Grained Reward Optimization: We propose an post-training alignment algorithm that fine-tunes the model to better align with fine-grained reward attributes.
@@ 2.1. Query Rewriting
The process of query rewriting, often termed query expansion or reformulation, is essential in enhancing e-commerce search technologies. It profoundly affects the shopping experience of users and the financial performance of e-commerce platforms. This approach can be generally classified into two classes: discriminative methods and generative methods.
Discriminative methods approach query rewriting as a retrieval task. Pseudo-relevance feedback methods (Xu and Croft, 2017b; Cao et al., 2008; Tao and Zhai, 2006; Xu and Croft, 2017a) enhance queries by identifying expansion terms from the top-ranked documents of an initial query, combining global corpus analysis with local feedback. Although these methods effectively address word mismatches, they are susceptible to semantic drift due to noisy or irrelevant top results. To address these challenges, researchers suggest utilizing a well-constructed thesaurus as a candidate set for query rewriting. (Bhogal et al., 2007; Mandal et al., 2019) However, they caution that the effectiveness of these methods is highly dependent on the thesaurus’s quality. If the thesaurus is inadequate, it may lead to query semantic drift, where the intended meaning of the query is compromised. Additionally, some approaches (Antonellis et al., 2008; Cui et al., 2002; Li et al., 2022) focus on generating candidate rewrites by leveraging search logs, using similar terms from users’ search histories as extensions. However, these search logs inherently display a bias towards popular queries due to the Matthew effect. Consequently, the training data derived from this method may fall short in optimizing for less frequently searched long-tail queries.
Generative methods, on the other hand, focus on the direct transformation of queries into keywords (Wang et al., 2024; Lee et al., 2018; Qi et al., 2020; Mohankumar et al., 2023). These methods involve transforming the input query directly into multiple bidwords, utilizing auto-regressive language models combined with trie-based decoding techniques to restrict the output space during inference. However, while previous methods ensure semantic relevance between input and output, they fall short in addressing value—a critical attribute in sponsored search advertising—which remains an open challenge. Our approach, as detailed in the following sections, effectively tackles this challenge by introducing innovative techniques that not only ensure semantic relevance but also optimize for value. By addressing this critical gap, our research has the potential to revolutionize sponsored search advertising, leading to more effective ad targeting and increased user satisfaction.
@@ 2.2. Preference Alignment
Reinforcement Learning from Human Feedback (RLHF) represents a pivotal technique for aligning LLMs with human preferences and values (Christiano et al., 2017; Bai et al., 2022; Ouyang et al., 2022), commonly applying the Bradley-Terry model (Bradley and Terry, 1952) to estimate the probability of a pairwise competition between two independently evaluated instances. The conventional RLHF pipeline is structured into three distinct phases: supervised fine-tuning (Zhou et al., 2024; Taori et al., 2023; Conover et al., 2023; Ding et al., 2023), reward model training (Gao et al., 2023; Chen et al., 2024; Havrilla et al., 2024b; Leike et al., 2018), and policy optimization (Schulman et al., 2017; Anthony et al., 2017). Notably, Proximal Policy Optimization (PPO) (Schulman et al., 2017) is extensively employed during the policy optimization phase. The RLHF framework has found broad applicability across various domains, including the mitigation of toxicity (Amini et al., 2024; Korbak et al., 2023), the assurance of safety (Dai et al., 2023), the enhancement of helpfulness (Wei et al., 2024) and the improvement of model reasoning capabilities (Havrilla et al., 2024a). Recent studies (Casper et al., 2023) have underscored the challenges inherent in the entire RLHF pipeline, spanning from the collection of preference data to the training of models. Furthermore, subsequent research has illuminated potential biases introduced by RLHF, such as the generation of verbose outputs (Dubois et al., 2024; Singhal et al., 2023; Wang et al., 2023).
@@ 3. Methodology
Traditional methods, such as standard beam search or top-k sampling, prioritize candidates with higher probabilities, often neglecting their value attributes. Although these approaches are effective in numerous contexts, they may fall short in applications where the value attributes are important. VALUE addresses this limitation by incorporating a Weighted Trie that allows for value-aware token selection. At each step of the decoding process, VALUE modifies the model’s output probabilities based on the value obtained from the trie, prioritizing tokens of higher value among those already highly relevant.
@@
3.1. System Overview
As can be seen from Figure 1, our entire system comprises the following key modules: ”LLM SFT and Alignment,” ”Weighted Trie Construction and Weight Momentum Update,” and ”Collaborative Inference.” For each generated token, we obtain the probability distribution from the LLM. Concurrently, we retrieve all possible candidates under the current prefix from the weighted trie along with their respective values. By integrating these values with the probability distribution, we derive a new probability distribution and sample the next token from it.

### v2

@@ Abstract.
Query-to-bidword (i.e., bidding keyword) rewriting is fundamental to sponsored search, transforming noisy user queries into semantically relevant and commercially valuable keywords. Recent advances in large language models (LLMs) improve semantic relevance through generative retrieval frameworks, but they rarely encode the commercial value of keywords. As a result, rewrites are often semantically correct yet economically suboptimal, and a reinforcement learning from human feedback (RLHF) stage is usually added after supervised fine-tuning (SFT) to mitigate this deficiency. However, conventional preference alignment frequently overemphasize the ordering of bidword values and is susceptible to overfitting, which degrades rewrite quality. In addition, bidword value changes rapidly, while existing generative methods do not respond to these fluctuations. To address this shortcoming, we introduce VALUE (Value-Aware Large language model for qUery rewriting via wEighted trie), a framework that integrates value awareness directly into generation and enhances value alignment during training. VALUE employs the Weighted Trie, a novel variant of the classical trie that stores real-time value signals for each token. During decoding, the framework adjusts the LLM’s token probabilities with these signals, constraining the search space and steering generation toward high-value rewrites. The alignment stage uses a fine-grained preference learning strategy that emphasizes stable, high-value differences and down-weights noisy or transient fluctuations, thereby improving robustness and reducing overfitting. Offline experiments show that VALUE significantly outperforms baselines in both semantic matching and value-centric metrics. Online A/B tests further revealed that our Revenue Per Mille (RPM) metric increased by 1.64%. VALUE has been deployed on our advertising system since October 2024 and served the Double Eleven promotions, the biggest shopping carnival in China.
@@
This author contributed to this work during his internship at Alimama.
@@ 1. Introduction
Sponsored search is a crucial element of modern search engines (Fain and Pedersen, 2006). Its effectiveness depends on accurately mapping user queries to relevant, high-value bidwords auctioned by advertisers. Query texts, however, are linguistically heterogeneous: they may contain imprecise phrasing, spelling errors, or synonyms. These variations obscure user intent and reduce matching accuracy. For instance, the query “drinks for fat person” should be reformulated as “low-sugar healthy drinks” before it can align with advertiser keywords. Meanwhile, the rapid expansion of e-commerce has further diversified query patterns. Current search behavior includes long-tail expressions, ambiguous product descriptions, and multi-intent requests, each linked to distinct advertisement sets with substantially different commercial value. Addressing these challenges requires query rewriting and keyword matching frameworks that incorporate both semantic equivalence and value awareness. Such systems must improve intent recognition while simultaneously optimizing for the expected value of candidate bidwords.
In recent years, generative query rewriting has garnered significant attention, with most state-of-the-art approaches following a three stage paradigm comprising supervised fine-tuning (SFT), value alignment, and Trie-constrained decoding (Peng et al., 2023; Chen et al., 2025; Wang et al., 2024a; Liu et al., 2025). In the SFT stage, domain-specific knowledge is injected into a large language model (LLM) through carefully curated rewriting corpora. The model then leverages its powerful semantic understanding, reasoning, and summarization capabilities to transform complex, noisy user queries into semantically canonical forms. During value alignment, reinforcement learning methods such as Proximal Policy Optimization (PPO), Direct Preference Optimization (DPO), Knowledge Transfer Optimization (KTO), and ORPO (Ouyang et al., 2022; Rafailov et al., 2024; Ethayarajh et al., 2024; Hong et al., 2024) adjust the model’s preferences so that the generated rewrites maximize not only semantic fidelity but also economic objectives including click-through rate and conversion likelihood. In the final generation stage, Trie-based decoding limits the output space to valid keyword candidates, ensuring syntactic correctness and retrieval feasibility. Historically, research has largely concentrated on better knowledge injection during SFT and on incorporating relevance, fluency, and factual consistency during alignment. Despite these advances, current systems remain inadequate in value awareness:
@@
    A fundamental misalignment exists between the commercial value ranking of keywords (driven by bid price and conversion potential) and their semantic relevance ranking in query rewriting. This divergence implies that semantically optimal rewrites rarely coincide with maximal-value keywords, creating inherent tension for generative rewriting systems. Recent frameworks like MoBGM (Liu et al., 2025) address this by jointly optimizing value and fidelity during alignment, yet require delicate balancing of training dynamics (e.g., reward weighting) to prevent objective dominance, which significantly increasing complexity.
@@
    Dynamic bid adjustments further cause real-time keyword value fluctuations. Consequently, value rankings observed during training become obsolete at inference, creating a temporal mismatch that undermines models relying on static/ historical signals. Current systems thus lack real-time value awareness, limiting adaptability to market dynamics and compromising economic optimality.
Figure 1. System Overview of our proposed VALUE Framework includes three parts: ”SFT and Alignment”, ”Weighted Trie Construction and Weight Momentum Update”, and ”Collaborative Inference”. During the LLM decoding process, we derive a value distribution from a Weighted Trie to constrain the generation space, and subsequently merge this with another distribution to produce the final output distribution.:
To address the aforementioned challenges, we propose shifting the focus of value integration from the training phase to the generation phase, where bidword value can be dynamically incorporated without sacrificing semantic relevance. We introduce a novel reward-guided inference framework, named as VALUE, and design a lightweight Weighted Trie structure that embeds bidword value directly into the decoding process, as illustrated in Figure 1. Specifically, during Trie construction, we annotate each node with downstream bidword value information, providing token-level commercial signals along the generation path. This allows the model to reason over both linguistic plausibility and economic return at each decoding step. To account for real-time bid fluctuations, we further develop a dynamic Trie update strategy that refreshes value annotations on the fly, ensuring the system remains responsive to changes in keyword valuation during inference. At each decoding step, the model queries the Weighted Trie for value-based rewards and modulates the token probability distribution accordingly. By maximizing cumulative output value during inference, this process decouples economic optimization from semantic learning, mitigating catastrophic forgetting and overfitting associated with value-centric training objectives. The framework therefore supports an explicit trade off between rewriting relevance and commercial utility.
To further bias generation toward consistently high-value bidwords, we refine DPO with adaptive loss weighting. The revised loss amplifies preference pairs exhibiting large, stable value gaps while down-weighting pairs subject to transient noise. This selectivity reduces sensitivity to unstable signals and improves generalization to long-tail and dynamically priced keywords. Overall, VALUE achieves a balanced integration of semantic fidelity and economic efficiency, bridging the gap between intent understanding and revenue maximization in generative query rewriting.
To evaluate the effectiveness of the proposed method, we conduct comprehensive offline experiments on historical online logs. Compared with the currently deployed production system, our approach achieves a 9.3% improvement in semantic relevance. Furthermore, large-scale online A/B testing demonstrates consistent gains in key business metrics: consumption increases by 1.53%, and revenue per thousand impressions (RPM) rises by 1.64%. Given these significant and sustained improvements, our method has been fully integrated into the live serving pipeline. Our contributions can be summarized as follows:
@@
    A pioneering reward-guided collaborative inference framework for query rewriting: We propose a novel generative rewriting framework that synergistically combines large language models (LLMs) with a value-augmented Weighted Trie, enabling dynamic, lightweight value awareness while preserving semantic fidelity, achieving a principled balance between relevance and commercial utility.
  2. (2)
    Innovative Trie structure with embedded value propagation: We redesign the classical Trie data structure by annotating each node with downstream bidword value information, providing token level commercial-value signals during generation. To our knowledge, this is the first work to transform the traditionally Trie-based constraint decoding into an active value-awareness mechanism within generative rewriting, allowing real-time adaptation to fluctuating bidword value through a dynamic update strategy.
  3. (3)
    Fine-grained reward optimization in post-training alignment: We introduce an enhanced preference learning algorithm that selectively emphasizes bidword pairs with large and stable value gaps, while down-weighting those susceptible to transient bid noise. This selective focus reduces noise exposure and mitigates overfitting risks, leading to more robust and generalizable value modeling.
@@ 2.1. Query Rewriting
The process of query rewriting, often termed query expansion or reformulation, is essential in enhancing e-commerce search technologies. This approach can be generally classified into two classes: discriminative methods and generative methods.
Discriminative methods frame query rewriting as a retrieval task. Pseudo-relevance feedback techniques (Xu and Croft, 2017a; Cao et al., 2008; Tao and Zhai, 2006; Xu and Croft, 2017b) expand queries by extracting terms from top-ranked initial documents, blending global corpus statistics with local feedback. While effective for lexical mismatch, these methods remain vulnerable to semantic drift from noisy top results. To mitigate this, curated thesauri have been proposed as candidate sources (Bhogal et al., 2007; Mandal et al., 2019), though their efficacy critically depends on thesaurus quality. Alternate approaches (Antonellis et al., 2008; Cui et al., 2002; Li et al., 2022) leverage search logs to generate rewrites, using historically similar terms as expansions. However, frequency bias in click data skews training toward head queries, yielding over-optimized embeddings that neglect long-tail queries. Dense retrieval models further struggle with distribution shifts, requiring costly mitigation strategies (e.g., meta-learning (Peng et al., 2023)) for long-tail generalization. Consequently, such methods inherently underperform for infrequent queries.
Generative methods, on the other hand, focus on the direct transformation of queries into bidwords (Wang et al., 2024b; Lee et al., 2018; Qi et al., 2020; Mohankumar et al., 2023; Wang et al., 2024a; Liu et al., 2025; Chen et al., 2025). Most of them adopt a standard pipeline of supervised fine-tuning, value alignment, and Trie-based decoding, which constrains outputs to bidwords that are both semantically relevant and commercially valuable. Earlier researches have improved supervised fine-tuning through prompt engineering to better exploit LLM reasoning and summarization, and has extended value alignment to incorporate objectives such as relevance and factual consistency. Nonetheless, existing systems still struggle to balance semantic fidelity with economic value and fail to adapt to real-time bidword fluctuations. Our approach, as detailed in the following sections, effectively tackles this challenge by introducing innovative techniques that not only ensure semantic relevance but also optimize for value.
@@ 2.2. Preference Alignment
Reinforcement Learning from Human Feedback (RLHF) aligns LLMs with human preferences by applying the Bradley–Terry model (Bradley and Terry, 1952) to estimate pairwise preferences (Christiano et al., 2017; Bai et al., 2022; Ouyang et al., 2022). A typical RLHF pipeline comprises three stages: supervised fine tuning (Zhou et al., 2024; Taori et al., 2023; Conover et al., 2023; Ding et al., 2023), reward model training (Gao et al., 2023; Chen et al., 2024; Havrilla et al., 2024b; Leike et al., 2018), and policy optimization (Schulman et al., 2017; Anthony et al., 2017). Proximal Policy Optimization (PPO) (Schulman et al., 2017) is widely used in the final stage. RLHF has been applied to reduce toxicity (Amini et al., 2024; Korbak et al., 2023), improve safety (Dai et al., 2023), enhance helpfulness (Wei et al., 2024), and strengthen reasoning (Havrilla et al., 2024a). However, challenges arise across the pipeline, from preference data collection to model training (Casper et al., 2023), and RLHF can introduce biases such as excessively long outputs (Dubois et al., 2024; Singhal et al., 2023; Wang et al., 2023).
Alternative alignment methods avoid explicit reward model training. Direct Preference Optimization (DPO) (Rafailov et al., 2024), Kahneman Tversky Optimization (KTO) (Ethayarajh et al., 2024), Simple Preference Optimization (SimPO) (Meng et al., 2024), and Monolithic Preference Optimization (ORPO) (Hong et al., 2024) optimize directly on preference pairs. SimPO and ORPO further remove the need for a reference model, simplifying training. Although these methods cannot sample preference pairs from an optimal policy, researchers augment data with outputs from a supervised fine-tuned model and employ iterative frameworks that periodically update the reference. Alignment as Reward Guided Search (ARGS) (Khanov et al., 2024) integrates alignment into decoding instead of training. Despite their advantages, these alignment techniques remain computationally intensive and must be retrained whenever value criteria change.
RLHF methods are commonly employed in generative query rewriting to infuse bidwords value information. However, a weak correlation between bidwords value ranking and relevance ranking often arises during rewriting. This misalignment leads to overfitting and degraded rewriting relevance in value-aware generative models. Our proposed VALUE framework addresses these issues through a dual-pronged approach: (1) leveraging the Weighted Trie for lightweight, real-time value awareness during generation, and (2) mitigating catastrophic forgetting of SFT-injected knowledge via optimized value preference alignment.
@@ 3. Methodology
As illustrated in Figure 1, VALUE operates in two primary phases: (1) a training phase integrating Supervised Fine-Tuning (SFT) with value preference alignment, and (2) a collaborative inference phase utilizing the Weighted Trie structure. During the SFT phase, preprocessed and relevant query-bidword pairs are employed to inject domain-specific rewriting knowledge for sponsored search into LLMs. Subsequently, we present our lightweight value-aware inference framework and the optimized value preference alignment method in detail.
We first introduce the construction of Weighted Trie and weight momentum update. Then, we will introduce collaborative inference based on Weighted Trie and the post-training alignment method we proposed in training phase. Finally, we will describe how VALUE is deployed for online serving.